# Dispersion on Time-Varying Graphs

### v1

@@ Abstract
The dispersion problem involves the coordination of $k\leq n$ agents on a graph of size $n$ to reach a configuration where at each node at most one agent can be present. It is a well-studied problem on static graphs. Also, this problem is studied on dynamic graphs with $n$ nodes where at each discrete time step the graph is a connected sub-graph of the complete graph $K_{n}$. An optimal algorithm is provided assuming global communication and 1-hop visibility of the agents. How this problem pans out on Time-Varying Graphs (TVG) is kept as an open question in the literature. In this work we study this problem on TVG where at each discrete time step the graph is a connected sub-graph of an underlying graph $G$ (known as a footprint) consisting of $n$ nodes. We have the following results even if only at most one edge from $G$ is missing in the connected sub-graph at any time step, and all agents start from a rooted initial configuration. Even with unlimited memory at each agent and 1-hop visibility, it is impossible to solve dispersion for $n$ co-located agents on a TVG in the local communication model. Furthermore, even with unlimited memory at each agent but without 1-hop visibility, it is impossible to achieve dispersion for $n$ co-located agents in the global communication model. From the positive side, the existing algorithm for dispersion on dynamic graphs with the assumptions of global communication and 1-hop visibility works on TVGs as well. This fact and the impossibility results push us to come up with a modified definition of the dispersion problem on TVGs, as one needs to start with more than $n$ agents if the objective is to drop the strong assumptions of global communication and 1-hop visibility. Then, we provide an algorithm to solve the modified dispersion problem on TVG starting with $n+1$ agents with $O(\log n)$ memory per agent while dropping both the assumptions of global communication and 1-hop visibility. This algorithm is optimal with respect to the number of agents (due to the impossibilities) and memory per agent.
@@ 1 Introduction
The dispersion problem, introduced in [2] on static graphs, involves the coordination of $k\leq n$ agents on the graph of size $n$ to reach a configuration with the constraint that at most one agent can be present at any node. This problem is generally applicable to real-world scenarios in which agents must coordinate and share resources located at various places. The goal is to minimize the total cost of solving the problem while ensuring that the cost of agents moving around on the graph is much lower than the cost of having more than one agent share a resource. The dispersion problem is also connected to three other problems: (i) scattering or uniform deployment of agents on a graph [4], (ii) exploration of agents on a graph [5], and (iii) load balancing on a graph [3]. The scattering problem involves coordinating and moving $k\leq n$ agents on a given graph with $n$ nodes so that they are uniformly distributed over the graph. When $k=n$, the scattering problem is equivalent to the dispersion problem. The problem of collaborative exploration requires agents to visit all the nodes of a given graph while coordinating with one another. This problem is related to dispersion as if $n$ agents need to disperse, they must explore the graph. Any solution to the dispersion problem can be applied to collaborative exploration as long as the assumptions and model parameters remain the same. Load balancing on graphs is a problem where resources need to be evenly distributed as much as possible among nodes on a graph in the least possible number of rounds. Typically, nodes control the allocation of resources, and there are constraints on how much load can pass through an edge in a round. This problem is similar to dispersion if we consider agents as resources.
@@
There are several works on dispersion [2, 7, 8, 14, 11, 15], and most of them deal with static graphs (graphs where the topology of the graph does not change with time). The addition of dynamism to a network makes the dispersion problem more challenging as well as practical. Agents may have to complete their tasks in environments that are also changing. In this work, we aim to solve the dispersion problem on dynamic graphs. Below we discuss the existing works on dispersion.
1.1 Related work
The problem of dispersion has been extensively studied in the literature [2, 7, 8, 14, 11, 15]. Augustine et al. in [2] introduced the dispersion problem. For the case where $k=n$, in their work [2], the authors present a memory lower bound of $\Omega(\log n)$ bits for each agent and a time lower bound of $\Omega(D)$ ($\Omega(n)$ on arbitrary graphs) for any deterministic algorithm on a given graph, where $D$ represents the diameter of the graph. In [2], authors have given two algorithms on arbitrary graphs for $k=n$: (i) algorithm with $O(m\cdot n)$ time using $O(\log n)$ bits at each agent and (ii) algorithm with $O(m)$ time using $O(n\cdot\log n)$ bits memory at each agent, where $m$ is the number of edges in the graph. Kshemkalyani et al. [7] provide an $\Omega(k)$ time lower bound for dispersion on arbitrary graphs for any $k\leq n$. In [7], authors have given three algorithms for arbitrary graphs for $k\leq n$: (i) $O(m)$ time using $O(k\cdot\log\Delta)$ memory at each agent, (ii) $O(\Delta^{D})$ time using $O(D\cdot\log\Delta)$ memory at each agent, and (iii) $O(m\cdot k)$ time using $O(\log(max\{k,\,\Delta\}))$ memory at each agent, where $\Delta$ is the maximum degree of the graph. Recently, Kshemkalyani et al. [8] provided an $O(\min\{m,\,k\cdot\Delta\}\cdot\log k)$-time with $O(\log n)$ memory per agent algorithm for arbitrary graphs, provided that the parameters $m$, $k$, $\Delta$ are known to agents. Kshemkalyani et al. [11] provided an algorithm that solves dispersion in $O(min\{m,\,k\cdot\Delta\})$ time with $\Theta(\log(k+\Delta))$ bits at each agent improving the time bound of the best previously known algorithm by $O(\log l)$ where $l\leq\frac{k}{2}$ and matching asymptotically the single-source depth-first search(DFS) traversal bounds [8]. Shintaku et al. [14] further present a dispersion algorithm that does not require any global knowledge and solves the dispersion problem on arbitrary graphs in $O(\min\{m,\,k\cdot\Delta\}\cdot\log l)$ rounds using $\Theta(\log(k+\Delta))$ bits of memory at each agent. Recently, Sudo et al. [15] presented even faster algorithms for both rooted and arbitrary configurations. For the rooted setting, they provide two algorithms. The first algorithm achieves dispersion in $O(k\cdot\log min\{k,\,\Delta\})$ time using $O(\log\Delta)$ bits of memory per agent. The second algorithm is time optimal, i.e., $O(k)$ time using $O(\Delta)$ memory per agent. The algorithm provided for the general setting requires $O(k\cdot\log k\cdot\log min\{k,\,\Delta\})$ using $O(\log(k+\Delta))$ memory per agent.
Recently, researchers in the field of distributed computing started studying dynamic graphs, where the topological changes are not sporadic or anomalous, but rather inherent to the nature of the network. Such a dynamic graph model was developed by Kuhn et al. [12] in 2010. Dispersion on the dynamic graphs is studied by Agarwalla et al. [1] and Kshemkalyani et al. [10]. In [1], the authors consider the following dynamic graph model on ring $\mathcal{R}$. In [1], the adversary can delete at most on edge from $\mathcal{R}$ at each round. Based on this dynamic graph model, they develop several deterministic algorithms for agents to achieve dispersion on a ring. In [10], the authors use a weaker dynamic graph model where the adversary can add or remove edges keeping the dynamic graph connected i.e., at any round, the dynamic graph $\mathcal{G}$ is connected. The authors show that it is impossible to solve dispersion on the dynamic graph $\mathcal{G}$ in the local communication model, even if $1$-hop visibility is available to the agents and each agent has unlimited memory. Also, it is impossible to solve dispersion on the dynamic graph $\mathcal{G}$ in the global communication model without $1$-hop visibility, even with unlimited memory at each agent. They provided an asymptotically optimal $\Theta(k)$ rounds algorithm in the global communication model with $1$-hop visibility.
The authors [10] ask the following question on the behavior of the dispersion problem on time-varying graphs (TVG) as TVG exhibits a stronger graph model than their model. It is easy to observe that the optimal algorithm of [10] also works for TVGs (as the graph model of TVG is stronger) in the global communication model with $1$-hop visibility. Due to the same reason, their impossibility results do not hold here on TVGs. In this work, we study the dispersion problem on TVG and provide impossibility results on the dispersion of $k\leq n$ agents in TVG, even from a rooted initial configuration, in the absence of either $1$-hop visibility or global communication. We further provide an algorithm to solve dispersion with $n+1$ agents from a rooted initial configuration on TVG dropping their assumptions of $1$-hop visibility and global communication. Below we provide our model description and the problem definition.
@@ 1.2 Model and the problem
Time-varying graph model: The system is represented as a time-varying graph (TVG), denoted by $\mathcal{G}=(V,E,\mathbb{T},\rho)$, where, $V$ represents the set of nodes, $E$ represents a set of edges, $\mathbb{T}$ represents the temporal domain, which is defined to be $\mathbb{Z}^{+}$ as in this model we consider discrete time steps, and $\rho:E\times\mathbb{T}\rightarrow\{0,1\}$ tells whether a particular edge is available at a given time. The graph $G=(V,\,E)$ is the underlying static graph of the dynamic graph $\mathcal{G}$, also termed as a footprint of $\mathcal{G}$. We denote $\delta_{v}$ as the degree of node $v$ in the footprint $G$, and $\Delta$ as the maximum degree in the footprint $G$. The graph $G$ is undirected and unweighted. The nodes of $G$ are anonymous (i.e., they have no IDs). The graph $G$ is port-labelled, i.e., each incident edge of a node $v$ is assigned a locally unique port number in $\{0,1,2,3,\cdots,\delta_{v}-1\}$. Specifically, each undirected edge $(u,\,v)$ has two labels, denoted as its port numbers at $u$ and $v$. Port numbering is local, i.e., there is no relationship between the port numbers at $u$ and $v$. We call a node as hole in round $t$ if there is no agent at that node in round $t$. We call a node a multinode in round $t$ if there are at least two agents present at that node in round $t$. We consider there is an adversary that can remove at most one edge from the footprint $G$ keeping $\mathcal{G}$ connected at each round.
Agent: We consider $k$ agents to be present initially at nodes of the graph $G$. Each agent has a unique identifier assigned from the range $[1,\,k]$. Each agent knows its ID. Agents do not have any prior knowledge of $k$, or any other graph parameters like $n$ and $\Delta$. The agents are equipped with $O(\log k)$ memory. An agent residing at a node (say $v$) in round $t$, knows $\delta_{v}$ in the footprint $G$ and all associated ports corresponding to node $v$ in $G$; however, the agent does not understand if any incident edge of node $v$ is missing in $\mathcal{G}$ in round $t$. To be more precise, agent $a$ being at node $v$ in round $t$ and $e_{v}$ being any edge associated with $v$, $a$ does not know the value of $\rho(e_{v},\,t)$.
@@
Visibility model: In this work, we use two types of visibility models: zero-hop visibility and $1$-hop visibility. In the zero-hop visibility model [2, 8, 9], an agent located at a particular node (say $v\in\mathcal{G}$) can see agents present at node $v$ in round $r$, and port numbers associated with node $v$. In $1$-hop visibility [1, 13], an agent (say $a_{i}$) located at a particular node (say $v\in\mathcal{G}$) can see any node (along with agent(s) present at that node, if any) that is the neighbours of $v$ in $\mathcal{G}$. If the adversary deletes an edge associated with a node $u$ with port $p$ at the start of a round $r$, the agent(s) present there cannot understand that an edge has been deleted in the zero-hop visibility model. However, if the agent tries to move through that edge in this round, it cannot make the move. In the 1-hop visibility model, if such an edge is deleted in a round, then the agents present there can see all the neighbours of $v$ that are connected to $v$ via the edges with the port number $\neq p$ at $v$ because according to the 1-hop visibility model, the agents can see the sub-graph induced by the nodes that are 1-hop away in this round. Therefore, agents can also recognize that no node is connected through port $p$, and therefore it can understand that the edge corresponding to port $p$ at $v$ is deleted at the start of this round.
Now we are ready to provide our problem definition.
@@
Dispersion on TVG: Let $\mathcal{G}$ be a TVG. Let $k$ agents be at the nodes of the graph $G$. At each round, the adversary can remove at most one edge from the footprint $G$ keeping $\mathcal{G}$ connected. If $k\leq n$, the agents reposition themselves such that each agent is at a distinct node of $G$; else, if $k>n$, the agents reposition themselves such that each node of $G$ contains at least one agent.
@@ 1.3 Our contribution
  • It is impossible to solve Dispersion on TVG with $k=n$ agents considering 1-hop visibility but without global communication, even if the agents start from a rooted configuration and have infinite memory (refer Section 2).
  • It is impossible to solve Dispersion on TVG with $k=n$ agents considering global communication but without 1-hop visibility, even if the agents start from a rooted configuration and have infinite memory (refer Section 2).
  • We present an algorithm (refer Section 3) that solves Dispersion on TVG with $n+1$ initially co-located agents in $O(n^{4}\cdot\log n)$ rounds that require $O(\log n)$ memory per agent considering local communication and zero visibility. This algorithm is optimal with respect to the number of agents given the impossibility results. Also, this algorithm is memory-optimal.
2 Impossibility Results
In this section, we show that it is impossible to solve dispersion deterministically in our model if either global communication or 1-hop visibility is omitted, even if agents start from a rooted configuration and have infinite memory. Recall that according to our model, the agents at a node $v$ cannot know whether an incident edge to node $v$ is missing or not.

### v2

# Ashish Saxena 11    Tanvir Kaur 11    Kaushik Mondal 11

@@ Abstract
Besides being studied over static graphs heavily, the dispersion problem is also studied on dynamic graphs with $n$ nodes where at each discrete time step the graph is a connected sub-graph of the complete graph $K_{n}$. An optimal algorithm is provided assuming global communication and 1-hop visibility of the agents. How this problem pans out on Time-Varying Graphs (TVG) is kept as an open question in the literature. In this work, we study this problem on TVG considering $k\geq 1$ agents where at each discrete time step the adversary can remove at most one edge keeping the underlying graph connected. We have the following main results considering all agents start from a rooted initial configuration. Global communication and 1-hop visibility are must to solve dispersion for $pn$ ($p\geq 1$) co-located agents on a TVG even if agents have unlimited memory and knowledge of $n$. We provide an algorithm that disperses $n+1$ agents on TVG by dropping both the assumptions of global communication and 1-hop visibility using $O(\log n)$ memory per agent. We extend this algorithm to solve dispersion with $pn+1$ agents with the same model assumptions.
@@ 1 Introduction
The dispersion problem, introduced in [3] on static graphs, involves the coordination of $k\leq n$ agents on the graph of size $n$ to reach a configuration with the constraint that at most one agent can be present at any node. This problem generally applies to real-world scenarios in which agents must coordinate and share resources at various places. The goal is to minimize the total cost of solving the problem while ensuring that the cost of agents moving around on the graph is much lower than the cost of having more than one agent share a resource. The dispersion problem is also connected to three other problems: (i) scattering or uniform deployment of agents on a graph [8], (ii) exploration of agents on a graph [9], and (iii) load balancing on a graph [6]. The scattering problem involves coordinating and moving $k\leq n$ agents on a given graph with $n$ nodes so that they are uniformly distributed over the graph. When $k=n$, the scattering problem is equivalent to the dispersion problem. The problem of collaborative exploration requires agents to visit all the nodes of a given graph while coordinating with one another. This problem is related to dispersion since if $n$ agents need to disperse, they must explore the graph. Load balancing on graphs is a problem where resources need to be evenly distributed as much as possible among nodes on a graph in the least possible number of rounds. Typically, nodes control the allocation of resources, and there are constraints on how much load can pass through an edge in a round. This problem is similar to dispersion if we consider agents as resources.
In [21], Molla et al. examine the dispersion problem for the first time for any value of $k\geq 1$. In this work, they have given the following definition of the dispersion.
Definition 1
Let $k$ agents be present at $n$ node graph $G$. Each node should have at most $\lceil k/n\rceil$ agents in the final configuration.
@@
There are several works on dispersion [3, 13, 14, 22, 17, 23, 21], and most of them deal with static graphs. The addition of dynamism to a network makes the dispersion problem more challenging as well as practical. Agents need to complete their tasks in environments that are also changing. In this work, we aim to solve the dispersion problem for any $k$ on dynamic graphs. Below, we discuss the model assumption and the problem definition.
@@ 1.1 Model and the problem
Time-varying graph model: The system is represented as a time-varying graph (TVG), denoted by $\mathcal{G}=(V,E,\mathbb{T},\rho)$, where, $V$ represents the set of nodes, $E$ represents a set of edges, $\mathbb{T}$ represents the temporal domain, which is defined to be $\mathbb{Z}^{+}$ as in this model we consider discrete time steps, and $\rho:E\times\mathbb{T}\rightarrow\{0,1\}$ tells whether a particular edge is available at a given time. The graph $G=(V,\,E)$ is the underlying static graph of the dynamic graph $\mathcal{G}$, also termed as a footprint of $\mathcal{G}$. We denote $\delta_{v}$ as the degree of node $v$ in the footprint $G$, and $\Delta$ as the maximum degree in the footprint $G$. The graph $G$ is undirected and unweighted. The nodes of $G$ are anonymous (i.e., they have no IDs). The graph $G$ is port-labelled, i.e., each incident edge of a node $v$ is assigned a locally unique port number in $\{0,1,2,3,\cdots,\delta_{v}-1\}$. Specifically, each undirected edge $(u,\,v)$ has two labels, denoted as its port numbers at $u$ and $v$. Port numbering is local, i.e., there is no relationship between the port numbers at $u$ and $v$. We consider there is an adversary that can remove edges from the footprint $G$, keeping $\mathcal{G}$ connected at each round.
Agent: We consider $k$ agents to be present initially at nodes of the graph $G$. Each agent has a unique identifier assigned from the range $[1,\,n^{c}]$, where $c$ is a constant. Each agent knows its ID. Agents are not aware of the values of $n$, $\Delta$, or $c$ unless these values are specifically mentioned. The agents are equipped with $O(\log n)$ memory. An agent residing at a node (say $v$) in round $t$ knows $\delta_{v}$ in the footprint $G$ and all associated ports corresponding to node $v$ in $G$; however, the agent does not understand if any incident edge of node $v$ is missing in $\mathcal{G}$ in round $t$. To be more precise, agent $a$ currently positioned at node $v$ at round $t$ does not know the value of $\rho(e_{v},t)$, where $e_{v}$ is an edge incident at node $v$. Such a model has been considered in [11, 12].
Visibility model: In this work, we use two types of visibility models: zero-hop visibility and $1$-hop visibility. In the zero-hop visibility model [3, 14, 15], an agent located at a particular node (say $v\in\mathcal{G}$) can see agents present at node $v$ in round $r$, and port numbers associated with node $v$. In $1$-hop visibility [1, 20], an agent (say $a_{i}$) located at a particular node (say $v\in\mathcal{G}$) can see any node (along with agent(s) present at that node, if any) that is the neighbours of $v$ in $\mathcal{G}$. If the adversary deletes an edge associated with a node $u$ with port $p$ at the start of a round $r$, the agent(s) present there cannot understand that an edge has been deleted in the zero-hop visibility model. However, if the agent tries to move through that edge in this round, it cannot make the move. In the 1-hop visibility model, if such an edge is deleted in a round, then the agents present there can see all the neighbours of $v$ that are connected to $v$ via the edges with the port number $\neq p$ at $v$ because according to the 1-hop visibility model, the agents can see the sub-graph induced by the nodes that are 1-hop away in this round. Therefore, agents can also recognize that no node is connected through port $p$, and therefore, agents can understand that the edge corresponding to port $p$ at $v$ is deleted at the start of this round.
@@
Now, we are ready to provide our problem definition.
Definition 2
@@
$k$-Dispersion on TVG: Let $\mathcal{G}$ be a TVG. Let $k$ $(\geq 1)$ agents be at the nodes of the footprint $G$. At each round, the adversary can remove at most one edge from $G$, keeping $\mathcal{G}$ connected. The agents need to reposition themselves such that each node contains at most $\lceil k/n\rceil$ agents.
In our problem definition, we consider the adversary can delete at most one edge from footprint $G$ at each round and this restriction is already considered in several existing works [4, 5, 12].
ID Range IC Dynamism $\bm{k}$ Necessary Assumptions Sufficient Assumptions Algorithm
[16] [1, k] Scattered Arbitrary edge addition or deletion maintaining the graph connected at each round $k\leq n$ 1-hop visibility, global comm. 1-hop visibility, global comm. $\Theta(k)$ rounds, $\Theta(\log k)$ memory per agent
This work $[1,\,n^{c}]$, $c\in\mathbb{N}$ Co-located 1-TVG $k=n+1$ 0-hop visibility, local comm. $O(n^{4}\cdot\log n)$ rounds, $\Theta(\log n)$ memory per agent
$k=pn+1$, $p>1$ 0-hop visibility, local comm., knowledge of $n$ $O(n^{4}\cdot\max(\log n,\,\log p))$ rounds, $\Theta(\log n)$ memory per agent
Co-located 1-TVG $k=pn$, $p\geq 1$ 1-hop visibility, global comm.
Scattered $f$-TVG $k\geq 1$ 1-hop visibility, global comm. $O(k)$ rounds, $\Theta(\log n)$ memory per agent (An extension of [16])
Table 1: This table shows the results of dispersion on different dynamic graph models. In this table, IC represents the initial configuration, $1$-TVG means the adversary can delete at most one edge per round from footprint $G$ while maintaining $\mathcal{G}$ connected, and $f$-TVG means the adversary can delete at most $f$ edges per round from footprint $G$ while maintaining $\mathcal{G}$ connected.
1.2 Related work
The problem of dispersion for $k$ $(\leq n)$ agents has been extensively studied on static graphs [3, 13, 14, 22, 17, 23]. Here we focus on the literature of dispersion on dynamic graphs only. Recently, researchers in the field of distributed computing started studying dynamic graphs, where the topological changes are not sporadic or anomalous but rather inherent to the nature of the network. Such a dynamic graph model was developed by Kuhn et al. [18] in 2010. Dispersion on the dynamic graphs for $k\leq n$ agents is studied by Agarwalla et al. [1] and Kshemkalyani et al. [16]. In [1], the authors consider the following dynamic graph model on ring $\mathcal{R}$. In [1], the adversary can delete at most one edge from $\mathcal{R}$ at each round. Based on this dynamic graph model, they develop several deterministic algorithms for agents to achieve dispersion on a ring. In [16], the authors use a weaker dynamic graph model than [1]. In [1], the adversary can delete at most one edge per round from the ring, but in [16], the adversary can add or remove edges, keeping the dynamic graph connected at each round. The authors show that it is impossible to solve dispersion on the dynamic graph in the local communication model, even if $1$-hop visibility is available to the agents and each agent has unlimited memory. Also, it is impossible to solve dispersion on the dynamic graph in the global communication model without $1$-hop visibility, even with unlimited memory at each agent. They provided an asymptotically optimal $\Theta(k)$ rounds algorithm in the global communication model with $1$-hop visibility.
In [2], Casteigts et al. defined the new dynamic graph paradigm, which is known as TVG. In this model, the exploration [11], gathering [19], and black-hole search problem [7, 4, 5, 12] have been studied. The basic difference between the dynamic graph model used in [16] and TVG is as follows. In [16], the footprint is a clique and for TVG’s the footprint can be graphs with less number of edges as well. Moreover, after edge deletion/addition, port numbers are adjusted in the model of [16], which is not the case in the TVG model. This allows us to understand missing edges in the presence of 1-hop visibility. That’s why TVG exhibits a stronger graph model than the dynamic graph model used in [16]. The authors [16] ask the following question on the behaviour of the dispersion problem on TVG. Since the graph model of TVG is stronger than the dynamic model used in [16], their impossibility results do not hold here on TVGs.
In this work, we study our $k$-Dispersion on TVG. In the preliminary version of this work, we demonstrated that 1-hop visibility and global communication are necessary for solving $n$-Dispersion on TVG, and we provided a high-level approach for addressing $(n+1)$-Dispersion on TVG. Here, we extend those results and provide our algorithms in detail.
@@ 1.3 Our contribution
  • Considering 1-hop visibility but without global communication, it is impossible to solve $k$-Dispersion on TVG when $k=pn$, $p\in\mathbb{N}$. This impossibility is valid even if the agents start from a rooted configuration, have knowledge of $n$ and are equipped with infinite memory (refer Section 2).
  • Considering global communication but without 1-hop visibility, it is impossible to solve $k$-Dispersion on TVG when $k=pn$, $p\in\mathbb{N}$. This impossibility is valid even if the agents start from a rooted configuration, have knowledge of $n$ and are equipped with infinite memory (refer Section 2).
  • We provide $\Omega(\log n)$ memory lower bound per agent to solve $k$-Dispersion on TVG (refer to Section 2), when $k>n$.
  • We present an algorithm that solves $(n+1)$-Dispersion on TVG when $n+1$ agents are initially co-located. This algorithm takes $O(n^{4}\cdot\log n)$ rounds and requires $O(\log n)$ memory per agent, considering local communication and zero visibility. This algorithm is memory-optimal (refer Section 3). If agents know $n$ and, $pn+1$ agents are co-located, then the same algorithm can be extended to solve $(pn+1)$-Dispersion considering local communication and zero visibility, where $p\in\mathbb{N}$ (refer to Section 4.1).
  • To solve $k$-Dispersion on TVG, we can use the existing algorithm of [16] when agents are equipped with 1-hop visibility and global communication. Further, the same algorithm of [16] with minor modifications works even when the adversary can delete any number of edges while keeping $\mathcal{G}$ connected in each round and $k(\geq 1)$ agents start from arbitrary initial configuration (refer to Section 4.2).
Figure 1: (a) If the adversary does not delete any edge at the beginning of round $r$, then this shows the computation of the agents. (b) If the adversary deletes $(u,\,u_{2})$ at the beginning of round $r$, then the view of agents at node $u_{1}$ does not change, but the view of agents at node $u$ and $u_{1}$ changes due to 1-hop visibility. In this case, agents at node $u$ can change $x_{1}$ to $x_{11}$, and agents at node $u_{2}$ can change $y_{3}$ to $y_{33}$ but agents at node $u_{1}$ do not change their decision because they can not understand deletion of edge $(u,\,u_{2})$ due to unavailability of global communication. Based on the cases mentioned in Theorem 2.1, the adversary decides which edge to delete at the start of round $r$. This outlines the high-level idea behind our proof Theorem 2.1.
2 Impossibility Results and Lower Bound
In this section, we show that it is impossible to solve $k$-Dispersion on TVG deterministically in our model if either global communication or 1-hop visibility is omitted, even if agents start from a rooted configuration. These impossibility results also hold if agents have infinite memory and knowledge of $n$. Recall that according to our model, the agents at a node $v$ cannot know whether an incident edge to node $v$ is missing or not unless agents are equipped with $1$-hop visibility.